# Multiscale Quantile Regression with Local Error Control

### v2

@@ A.2 Proof of Theorems 14
$\leq$ $\,\mathbb{P}\left(\max\,T_{I_{i}}(Z,\widehat{\theta}_{i})-q_{\alpha}(\lvert I_{i}\rvert)\geq 0,\forall\widehat{f}\in\Sigma\text{ with }\lvert T(\widehat{f})\rvert\leq K+k-1\right)$
$\leq$ $\,\mathbb{P}\left(\max\,T_{I_{i}}(Z,\widehat{\theta}_{i})-q_{\alpha}(\lvert I_{i}\rvert)\geq 0,\forall\widehat{f}\in\Sigma\text{ with }T(f)\subseteq T(\widehat{f})\text{ and }\lvert T(\widehat{f})\rvert\leq K+k-1\right)$
$\leq$ $\,\mathbb{P}\left(\max T_{I_{i}}(Z-f,\widehat{\theta}_{i}-f|_{I_{i}})-q_{\alpha}(\lvert I_{i}\rvert)\geq 0,\forall\widehat{f}\in\Sigma\text{ with }\lvert T(\widehat{f}-f)\rvert\leq k-1\right).$
Let $\widetilde{Z}=Z-f$, then each $\widetilde{Z}_{i}$ has $\beta$-quantile zero. The event in last lines implies that the estimator for zero function supplied by MUSCLE has at least $k$ change points. Thus, it follows from Proposition 8 that
$\mathbb{P}(\widehat{K}\geq K+k)\leq\mathbb{P}(T_{k}<n)\leq\alpha^{k}.$
@@
$\mathbb{P}(\widehat{K}_{n}\neq K_{n})\leq$  $\,4K_{n}\Bigg{(}\exp\left(-\left(\sqrt{n}\Omega_{n}-\frac{1}{\sqrt{2}}\underset{m\leq n}{\max}q_{\alpha}(n)-\sqrt{\log\left(\frac{2e}{\Lambda_{n}}\right)}\right)_{+}^{2}\right)+\exp\left(-n\Omega_{n}^{2}\right)\Bigg{)}+\alpha_{n}$
@@
where
$\Gamma_{1,n}$ $=\left(\sqrt{n}\Omega_{n}-\frac{1}{\sqrt{2}}\underset{m\leq n}{\max}q_{\alpha}(n)-\sqrt{\log\left(\frac{2e}{\Lambda_{n}}\right)}\right)_{+}^{2}-\log K_{n},$
$\Gamma_{2,n}$ $=n\Omega_{n}^{2}-\log K_{n}.$
Note that $\Gamma_{1,n}\leq\Gamma_{2,n}$ for all large $n$ and $\alpha_{n}\to 0$, it is sufficient to show that $\Gamma_{1,n}\to\infty$ as $n\to\infty$. In the first case: lim $\liminf\Lambda_{n}>0$ implies that both $\sqrt{\log(2e/\Lambda_{n})}$ and $K_{n}$ are finite, thus the condition
$\sqrt{n}\Omega_{n}-\frac{1}{\sqrt{2}}\underset{m\leq n}{\max}q_{\alpha}(n)\to\infty,$
ensures that $\Gamma_{1,n}\to\infty$.
For the second case: Note that $K_{n}\leq\Lambda_{n}^{-1}$ and $\sqrt{n}\Omega_{n}\geq c\sqrt{-\log\Lambda_{n}}$ for some $c>2$ and apply the basic inequality $\sqrt{x+y}-\sqrt{x}\geq y/(2\sqrt{x})$ for all $x,y\geq 0$, we then have
$\Gamma_{1,n}$ $\geq\left(c\sqrt{-\log\Lambda_{n}}-\frac{1}{\sqrt{2}}\underset{m\leq n}{\max}q_{\alpha}(n)-\sqrt{\log\left(\frac{2e}{\Lambda_{n}}\right)}\right)_{+}^{2}-\left(\sqrt{-\log\Lambda_{n}}\right)^{2}.$
$\geq\left((c-1)\sqrt{-\log\Lambda_{n}}-\frac{1}{\sqrt{2}}\underset{m\leq n}{\max}q_{\alpha}(n)-\frac{\log 2e}{2\sqrt{-\log\Lambda_{n}}}\right)_{+}^{2}-\left(\sqrt{-\log\Lambda_{n}}\right)^{2}$
$\geq\left((c-2)\sqrt{-\log\Lambda_{n}}-\frac{1}{\sqrt{2}}\underset{m\leq n}{\max}q_{\alpha}(n)-\frac{\log 2e}{2\sqrt{-\log\Lambda_{n}}}\right)_{+}^{2}\to\infty,$
where the last inequality comes from another basic inequality $(x+y)^{2}-x^{2}\geq y^{2}$ for all $x,y\geq 0$. ∎
@@
We consider the probability of $d_{H}(T_{n},\widehat{T}_{n})\geq\varepsilon_{n}$. Obviously,
$\mathbb{P}\left(d_{H}(T_{n},\widehat{T}_{n})\geq\varepsilon_{n}\right)\leq\mathbb{P}\left(\max_{k}\min_{j}\lvert\tau_{k}-\widehat{\tau}_{j}\rvert\geq\varepsilon_{n}\right)+\\ \mathbb{P}\left(\max_{k}\min_{j}\lvert\tau_{k}-\widehat{\tau}_{j}\rvert<\varepsilon_{n},\max_{j}\min_{k}\lvert\tau_{k}-\widehat{\tau}_{j}\rvert\geq\varepsilon_{n}\right).$ (23)
We bound both probabilities in (23) from above separately.
For the first probability: If $\max_{k}\min_{j}\lvert\tau_{k}-\widehat{\tau}_{j}\rvert\geq\varepsilon_{n}$, then there exists $k\in\{1,\dots,K_{n}\}$ such that on $I_{k}=(\tau_{k}-\varepsilon_{n},\tau_{k}+\varepsilon_{n})$ no change point is detected. Since $2\varepsilon_{n}\leq\Lambda_{n}$, the interval $I_{k}\subseteq(\tau_{k}-\Lambda_{n}/2,\tau_{k}+\Lambda_{n}/2)$. Applying the same technique as in the proof of Theorem 1 and replacing $\Lambda_{n}$ by $\varepsilon_{n}$ we obtain the first upper bound.
Next we study the second probability: We decompose $\widehat{T}_{n}$ into two disjoint sets $A_{n}$ and $B_{n}$, where
$A_{n}$ =$=\{\widehat{\tau}\in\widehat{T}_{n}:\min_{k}\lvert\widehat{\tau}-\tau_{k}\rvert<\varepsilon_{n}\},$
$B_{n}$ =$=\{\widehat{\tau}\in\widehat{T}_{n}:\min_{k}\lvert\widehat{\tau}-\tau_{k}\rvert\geq\varepsilon_{n}\}.$
Since $\max_{k}\min_{j}\lvert\tau_{k}-\widehat{\tau}_{j}\rvert<\varepsilon_{n}$, we have for any true change point $\tau_{k}$, there exists an estimated change point $\widehat{\tau}_{j(k)}$ such that $\lvert\tau_{k}-\widehat{\tau}_{j}\rvert<\varepsilon_{n}$, that is, $\widehat{\tau}_{j(k)}\in I_{k}$ for all $k=1,\dots,K_{n}$. Note that $I_{k}$’s are pairwise disjoint, thus $\widehat{\tau}_{j(k)}$’s are pairwise distinct. By definition, all $\widehat{\tau}_{j(k)}$’s belong to $A_{n}$ and therefore $\lvert A_{n}\rvert\geq K$. Furthermore, the event $\max_{j}\min_{k}\lvert\tau_{k}-\widehat{\tau}_{j}\rvert\geq\varepsilon_{n}$ implies that there exists a $j^{*}$ such that $\lvert\tau_{k}-\widehat{\tau}_{j^{*}}\rvert\geq\varepsilon_{n}$ for all $k=1,\dots,K_{n}$, i.e., $\widehat{\tau}_{j^{*}}\in B_{n}$ and $\lvert B_{n}\rvert\geq 1$. Since $\widehat{T}_{n}=A_{n}\dot{\sqcup}B_{n}$, we have $\lvert\widehat{T}_{n}\rvert\geq K_{n}+1$. Therefore, by Theorem 2 the second probability is upper bounded by $\mathbb{P}(\widehat{K}_{n}\geq K_{n}+1)\leq\alpha_{n}$. This finishes the proof. ∎
@@ A.3 Proof of Theorem 5
Let $\alpha\in(0,1)$. We first construct an upper bound for expectation of the number of false discoveries (denoted by $\rm{FD}(\alpha)$) given there is no true discovery.
@@
$\mathbb{E}(\rm{FD}(\alpha)|TD(\alpha)=0)\leq\frac{\alpha}{1-\alpha}\eqqcolon G(\alpha)$
@@
Note that it is sufficient to prove the claim for constant signal and we may assume that it equals to zero. We study the estimator $\widetilde{f}$ with $\widetilde{K}$ change points, which is defined in the proof of Proposition 8. Again, since MUSCLE minimises the number of estimated change points, the FD of MUSCLE is upper bounded by $\widetilde{K}$. Thus, apply Proposition 8 we have
$\mathbb{E}({\rm FD(\alpha)}|{\rm TD}(\alpha)=0)\leq\mathbb{E}(\widetilde{K})\leq\frac{\alpha}{1-\alpha}\eqqcolon G(\alpha).$
@@
Applying Lemma 11, Li et al., (2016, Lemma A.2) and the technique in Li et al., (2016, proof of Theorem 2.2) we have
@@
Next we prove the statement on OER. It follows from Theorem 2 that $\mathbb{P}(\widehat{K}-K\geq k)\leq\alpha^{k}$. Thus,
$\mathbb{E}((\widehat{K}-K)_{+})=\sum_{k=1}^{\infty}\mathbb{P}(\widehat{K}-K\geq k)\leq\frac{\alpha}{1-\alpha}.$
If $K=0$, then
@@
For $K\geq 1$,
$\frac{(\widehat{K}-K)_{+}}{\widehat{K}\vee 1}=\frac{(\widehat{K}-K)_{+}}{(\widehat{K}-K)_{+}+K}.$
Note that the function $x\mapsto x/(x+K)$ is strictly concave and monotone increasing on $[1,\infty)$. By Jensen’s inequality we have
$\mathbb{E}\left(\frac{(\widehat{K}-K)_{+}}{\widehat{K}+1}\right)\leq$ $\frac{\mathbb{E}((\widehat{K}-K)_{+})}{\mathbb{E}((\widehat{K}-K)_{+})+K}$
$\leq$ $\frac{\frac{\alpha}{1-\alpha}}{\frac{\alpha}{1-\alpha}+K}=\frac{\alpha}{(1-\alpha)K+\alpha}.$
Appendix B Additional materials of simulation study
Suppose that $Z_{1},\dots,Z_{n}$ are observations from Model 1 with underlying signal ${f}$. Let $\widehat{f}$ denote an estimate of ${f}$. Then the mean integrated squared error (MISE) of $\widehat{f}$ is given by
${\rm MISE}\left(\widehat{f}\right)=\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{f}(x_{i})-{f}(x_{i})\right)^{2}.$ (24)
Furthermore, we define the mean integrated absolute error (MIAE) of $\widehat{f}$ by
@@ Appendix B Additional materials of simulation study
Figure 9: Recovery of the blocks signal in (E3). In each panel, the overall performance over 200 repetitions is summarised as a boxplot and individual repetitions are jittered in dots with a low intensity. In the bottom left panel, the theoretical upper bound $\alpha=0.3$ on the FDR of MUSCLE is marked by a red dashed line. In the bottom right panel, the true numbers of change points in median and in distribution are $K=11$ (marked by a red dashed line) and $K=14$ (marked by a blue dashed line), respectively.
Figure 10: Recovery of the blocks signal in (E4). In each panel, the overall performance over 200 repetitions is summarised as a boxplot and individual repetitions are jittered in dots with a low intensity. In the bottom left panel, the theoretical upper bound $\alpha=0.3$ on the FDR of MUSCLE is marked by a red dashed line. In the bottom right panel, the true numbers of change points in median and in distribution are $K=11$ (marked by a red dashed line) and $K=14$ (marked by a blue dashed line), respectively.
Table 2: Statistical performance of procures mentioned in Section 5.1.1 in Gaussian (E1), scaled $t_{3}$ (E2), scaled Cauchy (E3), scaled centred $\chi_{3}^{2}$-distribution (E4) and mixed distribution (E5). Medians of each quantity and their mean absolute deviations (in brackets) are presented. For each quantity, the best performances are highlighted in bold.
Scenarios Methods MISE MIAE $L(T;\widehat{T})$ $d_{H}(T,\widehat{T})$ $\widehat{K}$ FDR V-meas. Time
@@
Scenarios Methods MISE MIAE $L(T;\widehat{T})$ $d_{H}(T,\widehat{T})$ $\widehat{K}$ FDR V-measure Time [s]

### v3

@@ A.2 Proofs of Theorems 14
$\leq\,$ $\mathbb{P}\Bigl(\max\,T_{I_{i}}(Z,\widehat{\theta}_{i})-q_{\alpha}(\lvert I_{i}\rvert)\geq 0,\forall\widehat{f}\in\Sigma\text{ with }\lvert T(\widehat{f})\rvert\leq K+k-1\Bigr)$
$\leq\,$ $\mathbb{P}\Bigl(\max\,T_{I_{i}}(Z,\widehat{\theta}_{i})-q_{\alpha}(\lvert I_{i}\rvert)\geq 0,\forall\widehat{f}\in\Sigma\text{ with }T(f)\subseteq T(\widehat{f})\text{ and }\lvert T(\widehat{f})\rvert\leq K+k-1\Bigr)$
$\leq\,$ $\,\mathbb{P}\Bigl(\max T_{I_{i}}(Z-f,\widehat{\theta}_{i}-f|_{I_{i}})-q_{\alpha}(\lvert I_{i}\rvert)\geq 0,\forall\widehat{f}\in\Sigma\text{ with }\lvert T(\widehat{f}-f)\rvert\leq k-1\Bigr).$
Let $\widetilde{Z}=Z-f$, and each $\widetilde{Z}_{i}$ has $\beta$-quantile zero. The event in last lines implies that the estimator for zero function by MUSCLE has at least $k$ change points. Thus, it follows from Proposition 8 that $\mathbb{P}(\widehat{K}\geq K+k)\leq\mathbb{P}(T_{k}<n)\leq\alpha^{k}.$
@@
$\mathbb{P}(\widehat{K}_{n}\neq K_{n})$
$\leq\,$ $4K_{n}\Bigg(\exp\left(-\left(\sqrt{n}\Omega_{n}-\frac{1}{\sqrt{2}}\underset{m\leq n}{\max}q_{\alpha}(n)-\sqrt{\log\frac{2e}{\Lambda_{n}}}\right)_{+}^{2}\right)+\exp\left(-n\Omega_{n}^{2}\right)\Bigg)+\alpha_{n}$
@@
where $\Gamma_{2,n}\coloneqq n\Omega_{n}^{2}-\log K_{n}$ and
$\Gamma_{1,n}\coloneqq\left(\sqrt{n}\Omega_{n}-\frac{1}{\sqrt{2}}\underset{m\leq n}{\max}q_{\alpha}(n)-\sqrt{\log\frac{2e}{\Lambda_{n}}}\right)_{+}^{2}-\log K_{n}.$
Note that $\Gamma_{1,n}\leq\Gamma_{2,n}$ for all large $n$ and $\alpha_{n}\to 0$. It is sufficient to show that $\Gamma_{1,n}\to\infty$ as $n\to\infty$. In the first case: $\liminf\Lambda_{n}>0$ implies that both $\sqrt{\log(2e/\Lambda_{n})}$ and $K_{n}$ are finite, and thus the condition $\sqrt{n}\Omega_{n}-\max_{m\leq n}q_{\alpha}(n)/{\sqrt{2}}\to\infty,$ ensures that $\Gamma_{1,n}\to\infty$.
For the second case: Note that $K_{n}\leq\Lambda_{n}^{-1}$ and $\sqrt{n}\Omega_{n}\geq c\sqrt{-\log\Lambda_{n}}$ for some $c>2$. Applying the basic inequality $\sqrt{x+y}-\sqrt{x}\geq y/(2\sqrt{x})$ for all $x,y\geq 0$, we then have
$\Gamma_{1,n}\geq\left(c\sqrt{-\log\Lambda_{n}}-\frac{\max_{m\leq n}q_{\alpha}(n)}{\sqrt{2}}-\sqrt{\log\frac{2e}{\Lambda_{n}}}\right)_{+}^{2}-\left(\sqrt{-\log\Lambda_{n}}\right)^{2}$
$\geq\,$ $\left((c-1)\sqrt{-\log\Lambda_{n}}-\frac{\max_{m\leq n}q_{\alpha}(n)}{\sqrt{2}}-\frac{\log 2e}{2\sqrt{-\log\Lambda_{n}}}\right)_{+}^{2}-\left(\sqrt{-\log\Lambda_{n}}\right)^{2}$
$\geq\,$ $\left((c-2)\sqrt{-\log\Lambda_{n}}-\frac{\max_{m\leq n}q_{\alpha}(n)}{\sqrt{2}}-\frac{\log 2e}{2\sqrt{-\log\Lambda_{n}}}\right)_{+}^{2}$
which goes to infinity. The last inequality above comes from the basic inequality $(x+y)^{2}-x^{2}\geq y^{2}$ for all $x,y\geq 0$. ∎
@@
Note that
$\mathbb{P}\left(d_{H}(T_{n},\widehat{T}_{n})\geq\varepsilon_{n}\right)\leq\mathbb{P}\left(\max_{k}\min_{j}\lvert\tau_{k}-\widehat{\tau}_{j}\rvert\geq\varepsilon_{n}\right)\\ +\mathbb{P}\left(\max_{k}\min_{j}\lvert\tau_{k}-\widehat{\tau}_{j}\rvert<\varepsilon_{n},\max_{j}\min_{k}\lvert\tau_{k}-\widehat{\tau}_{j}\rvert\geq\varepsilon_{n}\right).$ (22)
For the first probability, if $\max_{k}\min_{j}\lvert\tau_{k}-\widehat{\tau}_{j}\rvert\geq\varepsilon_{n}$, then there exists a $k\in\{1,\dots,K_{n}\}$ such that on $I_{k}=(\tau_{k}-\varepsilon_{n},\tau_{k}+\varepsilon_{n})$ no change point is detected. Since $2\varepsilon_{n}\leq\Lambda_{n}$, we have the interval $I_{k}\subseteq(\tau_{k}-\Lambda_{n}/2,\tau_{k}+\Lambda_{n}/2)$. Applying the same technique as in the proof of Theorem 1 and replacing $\Lambda_{n}$ by $\varepsilon_{n}$, we obtain the first upper bound.
For the second probability, we decompose $\widehat{T}_{n}$ into two disjoint sets $A_{n}=\{\widehat{\tau}\in\widehat{T}_{n}:\min_{k}\lvert\widehat{\tau}-\tau_{k}\rvert<\varepsilon_{n}\}$ and $B_{n}=\{\widehat{\tau}\in\widehat{T}_{n}:\min_{k}\lvert\widehat{\tau}-\tau_{k}\rvert\geq\varepsilon_{n}\}.$ Since $\max_{k}\min_{j}\lvert\tau_{k}-\widehat{\tau}_{j}\rvert<\varepsilon_{n}$, we have for any true change point $\tau_{k}$, there exists an estimated change point $\widehat{\tau}_{j(k)}$ such that $\lvert\tau_{k}-\widehat{\tau}_{j}\rvert<\varepsilon_{n}$, that is, $\widehat{\tau}_{j(k)}\in I_{k}$ for all $k=1,\dots,K_{n}$. Note that $I_{k}$’s are pairwise disjoint. Thus $\widehat{\tau}_{j(k)}$’s are pairwise distinct. By definition, all $\widehat{\tau}_{j(k)}$’s belong to $A_{n}$ and therefore $\lvert A_{n}\rvert\geq K$. Furthermore, the event $\max_{j}\min_{k}\lvert\tau_{k}-\widehat{\tau}_{j}\rvert\geq\varepsilon_{n}$ implies that there exists a $j^{*}$ such that $\lvert\tau_{k}-\widehat{\tau}_{j^{*}}\rvert\geq\varepsilon_{n}$ for all $k=1,\dots,K_{n}$, i.e., $\widehat{\tau}_{j^{*}}\in B_{n}$ and $\lvert B_{n}\rvert\geq 1$. Since $\widehat{T}_{n}=A_{n}\sqcup B_{n}$, we have $\lvert\widehat{T}_{n}\rvert\geq K_{n}+1$. Therefore, by Theorem 2, the second probability is upper bounded by $\mathbb{P}(\widehat{K}_{n}\geq K_{n}+1)\leq\alpha_{n}$. ∎
@@ A.3 Proof of Theorem 5
Let $\alpha\in(0,1)$. We first construct an upper bound for expectation of the number of false discoveries $\rm{FD}(\alpha)$ given there is no true discovery.
@@
$\mathbb{E}(\rm{FD}(\alpha)\mid TD(\alpha)=0)\leq\frac{\alpha}{1-\alpha}\eqqcolon G(\alpha)$
@@
Note that it is sufficient to prove the claim for constant signal and we may assume that it equals to zero. We study the estimator $\widetilde{f}$ with $\widetilde{K}$ change points, which is defined in the proof of Proposition 8. Again, since MUSCLE minimizes the number of estimated change points, the FD of MUSCLE is upper bounded by $\widetilde{K}$. Thus, applying Proposition 8, we have $\mathbb{E}({\rm FD(\alpha)}\mid{\rm TD}(\alpha)=0)\leq\mathbb{E}(\widetilde{K})\leq\frac{\alpha}{1-\alpha}\eqqcolon G(\alpha).$
@@
Applying Lemma 11, Li et al., (2016, Lemma A.2) and the proof technique in Li et al., (2016, Theorem 2.2), we have
@@
Next we prove the statement on OER. It follows from Theorem 2 that $\mathbb{P}(\widehat{K}-K\geq k)\leq\alpha^{k}$. Thus, $\mathbb{E}((\widehat{K}-K)_{+})=\sum_{k=1}^{\infty}\mathbb{P}(\widehat{K}-K\geq k)\leq{\alpha}/({1-\alpha}).$ If $K=0$, then
@@
For $K\geq 1$, $\frac{(\widehat{K}-K)_{+}}{\widehat{K}\vee 1}=\frac{(\widehat{K}-K)_{+}}{(\widehat{K}-K)_{+}+K}.$ The function $x\mapsto x/(x+K)$ is strictly concave and monotone increasing on $1,\infty)$. By Jensen’s inequality
$\mathbb{E}\left(\frac{(\widehat{K}-K)_{+}}{\widehat{K}+1}\right)\leq\frac{\mathbb{E}((\widehat{K}-K)_{+})}{\mathbb{E}((\widehat{K}-K)_{+})+K}\leq\frac{\frac{\alpha}{1-\alpha}}{\frac{\alpha}{1-\alpha}+K}=\frac{\alpha}{(1-\alpha)K+\alpha}.\qed$
Appendix B Additional Materials for Section 5
Suppose that $Z_{1},\dots,Z_{n}$ are observations from Model 1 with underlying signal f{f}. Let $\widehat{f}$ denote an estimate of f{f}. Then the mean integrated squared error (MISE) of $\widehat{f}$ is given by
${\rm MISE}\left(\widehat{f}\right)=\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{f}(x_{i})-{f}(x_{i})\right)^{2},$ (23)
and the mean integrated absolute error (MIAE) of $\widehat{f}$ by
@@ Appendix B Additional Materials for Section 5
Figures 9, 10 and 2 present additional results for Section 5.1.1, while Figure 11 corresponds to Section 5.2.1.
Figure 9: Recovery of the blocks signal in (E3). In each panel, the overall performance over 200 repetitions is summarized as a boxplot and individual repetitions are jittered in dots with a low intensity. In the bottom left panel, the theoretical upper bound $\alpha=0.3$ on the FDR of MUSCLE is marked by a red dashed line. In the bottom right panel, the true numbers of change points in median and in distribution are $K=11$ (marked by a red dashed line) and $K=14$ (marked by a blue dashed line), respectively.
Figure 10: Recovery of the blocks signal in (E4). In each panel, the overall performance over 200 repetitions is summarized as a boxplot and individual repetitions are jittered in dots with a low intensity. In the bottom left panel, the theoretical upper bound $\alpha=0.3$ on the FDR of MUSCLE is marked by a red dashed line. In the bottom right panel, the true numbers of change points in median and in distribution are $K=11$ (marked by a red dashed line) and $K=14$ (marked by a blue dashed line), respectively.
Table 2: Statistical performance of procedures in Section 5.1.1 in Gaussian (E1), scaled $t_{3}$ (E2), scaled Cauchy (E3), scaled and centered $\chi_{3}^{2}$-distribution (E4) and mixed distribution (E5). Medians of each criterion and their mean absolute deviations (in brackets) are presented, with the best performances highlighted in bold.
Scenario Method MISE MIAE L(T;T^)L($\widehat{T}$$)d_{H}(T,\widehat{T}$) $\widehat{K}$ FDR V-measure Time [s]
@@
Table 3: Continuation of Table 2.
Scenario Method MISE MIAE L(T;T^)L($\widehat{T}$$)d_{H}(T,\widehat{T}$) $\widehat{K}$ FDR V-measure Time [s]